OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through (1, 2) and is parallel to the line whose equation is 3x - 4y + 12 = 0? 3x + 4y - 5 = 0 , 3x - 4y + 5 = 0 , 4x + 3y + 5 = 0

OpenStudy (mertsj):

Parallel lines have equal slopes. The slope of the given line is 3/4. Now plug into point-slope form. y-2=3/4(x-1) y-2=3/4x -3/4 y=3/4x+5/4